# Share the Sugar

### v1

@@ Abstract
We provide a general argument against value incomparability, based on a new style of impossibility result. In particular, we show that, against plausible background assumptions, value incomparability creates an incompatibility between two very plausible principles for ranking lotteries: a weak “negative dominance” principle (to the effect that Lottery 1 can be better than Lottery 2 only if some possible outcome of Lottery 1 is better than some possible outcome of Lottery 2) and a weak form of ex ante Pareto (to the effect that, if Lottery 1 gives an unambiguously better prospect to some individuals than Lottery 2, and equally good prospects to everyone else, then Lottery 1 is better than Lottery 2). After spelling out our results, and the arguments based on them, we consider which principle the proponent of incomparability ought to reject.
Keywords: value theory, population ethics, ethics & decision theory, incompleteness, dominance, Pareto, opaque sweetening
@@ 1 Introduction
It often seems, when comparing two possible outcomes, that neither is better than the other, nor are they exactly equally good. Suppose that Patricia is deciding whether to pursue a career as an artist or a banker. Being an artist would give her an unlimited creative outlet and the freedom of being her own employer. Being a banker would give her financial stability and regular opportunities for international travel. Even setting aside her many uncertainties and imagining maximally specific ways that each of these lives might go, it seems that neither life would be better for Patricia overall—and that they would not be exactly equal, in the sense that a slight but noticeable improvement to one life (say, adding \$1,000 in salary) would not be enough to make it better than the other. Similarly, from an impersonal point of view (“the point of view of the universe”, or of social welfare), the outcome where Patricia becomes an artist could be neither better nor worse nor exactly as good as the outcome where she becomes a banker. [^1] Philosophers who have argued that lives can be incomparable in this way include Raz (1985, 1986), Chang (2002), and Hedden and Muñoz (2023), among many others.
@@
Separate from this sort of incomparability between lives, many philosophers have claimed that a structurally similar phenomenon arises when comparing outcomes in which different individuals exist. Some have claimed that, from a personal point of view, existence is always incomparable with non-existence: no matter how good or bad an individual’s life, we can never say that it is better, worse, or exactly as good for her as non-existence (Broome, 1999, p. 168). Some have claimed that the same is true from an impersonal point of view: two possible worlds or outcomes that are identical except that some individual exists in one but not in the other are incomparable in terms of overall value, whatever life that individual lives in the world where they exist (Bader, 2022, 2023). Others have claimed that this is true only of some lives: a life with the right mix of triumphs and tragedies may be incomparable with non-existence, from an impersonal point of view (Gustafsson, 2020; Thornley, 2022).
@@
This paper presents a new argument against these and other forms of incomparability. Suppose that two possible lives, like artist and banker, are indeed incomparable. Let $a$ denote a fully specified artist’s life and $b$ a fully specified banker’s life, assumed to be incomparable, with $a^{+}$ a slightly improved version of the artist’s life that is still incomparable with $b$. Now suppose we face a choice between the following two lotteries, based on the same fair coin flip, where our choice will affect only two individuals, $P_{1}$ and $P_{2}$:
@@
There is, on the one hand, a strong argument that neither lottery is better than the other. First, if $a$ and $b$ are incomparable from a personal point of view, then it seems that two outcomes that differ only in that some individual enjoys life $a$ in one and life $b$ in the other should also be incomparable. Thus, $(a,b)$ and $(b,a)$ are both incomparable with $(a,a)$ and $(b,b)$, where the first and second terms in each ordered pair represent the lives of $P_{1}$ and $P_{2}$ respectively. Moreover, it is part and parcel of incomparability that it is preserved by small enough improvements, so $(a^{+},a^{+})$, which is at most slightly better than $(a,a)$, should likewise be incomparable with $(a,b)$ and $(b,a)$. We then find that every possible outcome of Lottery 1 (the Heads outcome and the Tails outcome) is incomparable with every possible outcome of Lottery 2. This suggests that neither lottery is strictly better than the other.
@@
On the other hand, there is also a strong argument that Lottery 2 is better than Lottery 1: Lottery 1 gives both $P_{1}$ and $P_{2}$ a 50% chance of life $a$ and a 50% chance of life $b$. Lottery 2 gives each of them a 50% chance of life $a^{+}$ and a 50% chance of life $b$. Since $a^{+}$ is better than $a$, Lottery 2 seems clearly more desirable from the point of view of each affected individual, and therefore better overall.
@@
Taken together, these arguments amount to an apparent reductio of the supposition that $a$ and $b$ are incomparable. Similar arguments, as we will see, can be mounted against many other forms of incomparability in the ranking of social outcomes: between larger and smaller populations, between competing desiderata like total welfare, average welfare, and equality, between benefits conferred on different infinite populations, and more. The result will be a new, general argument for completeness of betterness, turning on considerations of a quite different character than the recent linguistic arguments of Dorr et al. (2021, 2023). The central premises of this reductio are two dominance-like principles: first, that if no possible outcome of lottery $L_{i}$ is better than any possible outcome of lottery $L_{j}$, then $L_{i}$ is not better than $L_{j}$; and second, that if $L_{i}$ is better for everyone affected than $L_{j}$ (in a very strict sense, to be explained), then it is better overall. We will call the first principle Negative Dominance and the second principle Personal Good.
@@
In §2, we formalize and motivate Negative Dominance, comparing it to the significantly stronger principle of “Recognition”, which plays a central role in Caspar Hare’s “opaque sweetening” puzzle (Hare, 2010). In §3, we similarly motivate and explain Personal Good. In §4, we offer a more exact version of the above argument, showing that, against modest background assumptions, Negative Dominance and Personal Good rule out incompleteness in the personal ranking of lives, and by extension those forms of incompleteness in the impersonal ranking of outcomes that derive from incomparability between lives. We show how this argument differs from a prima facie related one (though developed for a different purpose) due to Nebel (2020). §5 then presents a more general argument against incompleteness in the impersonal ranking of outcomes, likewise based on Negative Dominance and Personal Good. §6 takes the argument a step further by showing that Personal Good plus plausible auxiliary assumptions allows the ranking of a large class of lotteries to be reduced to the ranking of outcomes, so that if outcomes are completely ordered by impersonal value, this class of lotteries will be too. Finally, in §7, we consider whether either Negative Dominance or Personal Good can be plausibly rejected. While neither principle is sacrosanct, we think that the most natural move for those committed to incomparability is to reject Negative Dominance. This move is especially natural for those who understand evaluative incompleteness as indeterminacy between complete rankings. It has the cost, however, of ruling out the “deferentialist” position that has been endorsed by several prominent fans of incompleteness (Schoenfield, 2014; Bales et al., 2014; Doody, 2019a, b, 2021, 2023).
@@ 2 Negative Dominance
We will be interested in the implications of certain decision-theoretic principles in an ethical context, so we begin with an abstract decision-theoretic framework which we will subsequently give an ethical interpretation. At the most abstract level, we are interested in rankings of lotteries over outcomes. Here, outcomes can be understood as complete possible worlds, or as equivalence classes of equally good worlds, or as specifications of all evaluatively significant features of a world. A lottery is any probability distribution over outcomes. We will say that an outcome is in a lottery if the lottery assigns it positive probability. Except for one brief stretch in §7, we will limit our attention to lotteries with only finitely many possible outcomes in them. Finally, we will assume that lotteries can be compared in terms of impersonal betterness. Specifically, we will take as primitive an at least as impersonally good as relation on lotteries. If lottery $L$ is at least as good as lottery $L^{\prime}$ but $L^{\prime}$ is not at least as good as $L$, we say that $L$ is strictly impersonally better than $L^{\prime}$. If they are each at least as good as each other, we say that they are equally impersonally good. If neither relation holds, we say that they are impersonally incomparable. The impersonal betterness relation on lotteries induces an impersonal betterness relation on outcomes: If the lottery that yields $o$ with certainty is at least as good as the lottery that yields $o^{\prime}$ with certainty, we say that $o$ is at least as impersonally good as $o^{\prime}$ (and likewise for other evaluative relations).
@@
Our interest here will be in the ethics of population and distribution. In this context, we assume that the evaluatively significant features of an outcome are (i) which individuals exist in that outcome and (ii) how their lives go. Thus, we will identify outcomes with specifications of only these features. In particular, we will assume a countable set of possible individuals (individuals who exist in some possible world/outcome) and a set of possible lives (complete specifications of everything that matters, morally or prudentially, about an individual’s life—e.g., the totality of their experiences). Importantly, the set of possible “lives” includes non-existence (“the empty life”). Given this background, we will understand outcomes as functions from the set of possible individuals to the set of possible lives, understood as specifying which individuals exist in a given outcome and what lives they lead. We will sometimes call such outcomes populations, when we want to emphasize the fact that they have less information in them than a possible world, but in general we will use the terms outcome and population interchangeably, and choose one or the other depending on which feature of these objects we wish to emphasize. Finally, alongside the impersonal betterness relation introduced above, we will assume a personal betterness relation on lives. [^2] We could define this relation also on lotteries over lives, but our principles and arguments will only make reference to personal betterness on lives, so we’ll make do with that more minimal relation. As above, we take as primitive an at least as (personally) good as relation on lives. If life $a$ is at least as good as life $b$ but $b$ is not at least as good as $a$, we say that $a$ is strictly (personally) better than $b$. If they are each at least as good as each other, we say that they are equally (personally) good. If neither relation holds, we say that they are (personally) incomparable. (From now on, we will usually omit “personal” and “impersonal” where it is clear from context which relation is meant.) [^3] This setup abstracts from a number of features which will not be important below. Most obviously we will be assuming that (the evaluatively significant aspects of) a life which is enjoyed by one individual could also be enjoyed by another. One might contest this on many grounds: perhaps, for instance, the particularity of friendships or familial relations might be thought to be evaluatively significant although an individual cannot be a friend of or a parent to themselves. A more perspicuous setting might take as primitive a set of worlds and individuals, and an “at least as (personally) good as” ranking of (lotteries over) relevant centered worlds, i.e. pairs of worlds and individuals. “Lives” could then be understood as equivalence classes of centered worlds with respect to the relation of exactly equally as (personally) good as. Populations/outcomes in our sense could be seen as equivalence classes of worlds under the equivalence relation which takes $w$ and $w^{\prime}$ to be equivalent if and only if for all $i$, $(w,i)$ is equally personally good as $(w^{\prime},i)$. Everything we do below could be done in a more specific setting like this one, but the extra detail would needlessly clutter the exposition.
@@
A betterness relation is complete if no two entities in its domain are incomparable—that is, for any two entities, either one is strictly better than the other, or they are equally good. Our interest in this paper will be in theories on which the impersonal betterness relation is incomplete. At some points we will also consider theories which hold that the personal betterness relation is incomplete as well.
@@
If for every outcome $o_{1}$ in lottery $L_{1}$ and every outcome $o_{2}$ in lottery $L_{2}$, $o_{1}$ is not at least as good as $o_{2}$, then $L_{1}$ is not strictly better than $L_{2}$.
We can argue for this principle as follows. Suppose you are considering a lottery with objective chances, that you know these chances, and that your preferences coincide with the objective relation of betterness over the outcomes of those lotteries. Then consider the principle:
If no epistemically possible outcome of $L_{1}$ is preferable to any epistemically possible outcome of $L_{2}$, then you are not required to choose $L_{1}$ over $L_{2}$.
The idea here is that, given the betterness facts about the outcomes (reflected in your preferences), there can’t be a reason, justification, or rationale for choosing $L_{1}$ that would be stronger than or defeat your reasons, justifications, or rationales for choosing $L_{2}$. [^4] For related discussion, see Hare (2022, Ch. 4). Given that there is no reason in favor of $L_{1}$ that outweighs or defeats the reasons in favor of $L_{2}$, it can’t be that you are required to choose $L_{1}$. But this means that $L_{1}$ can’t be strictly better. If it were, then since we assumed that you knew all the betterness facts, and had preferences in line with them, you would be required to choose it. So it must be that $L_{1}$ is not strictly better in this case.
In his important paper “Take the Sugar”, Caspar Hare proposes a principle, which he calls “Recognition”, that shares some of the appeal of Negative Dominance, but (as we will now argue) is too strong in key cases. [^5] Our title is a play on Hare’s, since we discuss a puzzle and a principle closely related to his, but in a context of multi-patient moral decision-making rather than individual prudential decision-making. Seeing the contrast between these two principles will bolster the case for Negative Dominance.
@@
Hare works in a slightly different setting than ours, with two different primitive objects: states and prizes. In this setting, probabilities are defined on states, while betterness is defined in the first instance on prizes. Gambles are functions from states to prizes. If we think of outcomes as state-prize pairs, then gambles induce lotteries, since a probability distribution on states yields, via the gamble, a probability distribution on prizes-cum-states. But gambles have more structure than lotteries, allowing us to contrast a possible outcome of a given gamble with what would have happened (in the same state) if a different gamble had been chosen. [^6] For those familiar with the distinction, Hare’s framework is in this respect akin to Savage’s, with the objects of choice understood as functions from states to prizes, while our framework is akin to von Neumann & Morgenstern’s.
@@
If for every state $s$, the prize that gamble $L_{1}$ yields in $s$ is not at least as good as the prize that gamble $L_{2}$ yields in $s$, then $L_{1}$ is not strictly better than $L_{2}$.
If we understand outcomes as state-prize pairs, as above, then Statewise Negative Dominance is strictly stronger than Negative Dominance, so anyone who accepts the former (e.g. Schoenfield (2014); Bales et al. (2014); Doody (2023)) must also accept the latter.
But even those who reject Statewise Negative Dominance (like Hare himself, Bader (2018), and Rabinowicz (2021)) still have good reason to endorse Negative Dominance. In particular, as Bader (2018) has emphasized, Statewise Negative Dominance conflicts with another, more widely accepted dominance principle:777 Manzini and Mariotti (2008, 310) earlier independently presented a very similar argument for an incompatibility between Stochastic Dominance (their “Preference Sure Thing”) and their “Vague Sure Thing Principle” (VST). Their VST is even stronger than Statewise Negative Dominance, but the line of thought which leads to incompatibility is the same.
@@
(i) If for every outcome $o$, $L_{1}$ has at least as great a probability as $L_{2}$ of yielding an outcome at least as impersonally good as $o$, then $L_{1}$ is at least as impersonally good as $o$. (ii) If in addition, for some $o$, $L_{1}$ has a strictly greater probability of yielding an outcome at least as impersonally good as $o$, then $L_{1}$ is strictly impersonally better than $L_{2}$. [^8] Russell (forthcoming) shows that when the betterness ordering is incomplete (and in other, more outré contexts), this standard formulation of Stochastic Dominance gives the wrong results and must be revised. But the difference between the standard formulation and Russell’s proposed revision won’t matter here.
@@
When condition (i) is satisfied, we say that $L_{1}$ weakly stochastically dominates $L_{2}$. When condition (ii) is also satisfies, we say that $L_{1}$ strictly stochastically dominates (or simply stochastically dominates) $L_{2}$.

### v2

# A Dominance Argument Against Incompleteness

@@ Abstract
This article presents a new argument against many forms of moral and prudential value incompleteness. The argument relies on two central principles: (i) a weak “negative dominance” principle, to the effect that Lottery 1 is better than Lottery 2 only if some possible outcome of Lottery 1 is better than some possible outcome of Lottery 2, and (ii) a weak form of ex ante Pareto, to the effect that, if Lottery 1 gives an unambiguously better (stochastically dominant) prospect to some individuals than Lottery 2, and equally good prospects to everyone else, then Lottery 1 is better than Lottery 2. Given modest auxiliary assumptions, these two principles rule out incompleteness in the prudential ranking of individual lives, and many forms of incompleteness in the moral rankings of outcomes and lotteries.
Forthcoming in Philosophical Review.
Keywords value theory, population ethics, incompleteness, incomparability, dominance, Pareto, opaque sweetening
@@ 1 Introduction
It often seems, when comparing two alternatives, that neither is better than the other, nor are they equally good. Suppose that Patricia is deciding whether to pursue a career as an artist or a banker. Being an artist would give her an unlimited creative outlet and the freedom of being her own employer. Being a banker would give her financial stability and regular opportunities for international travel. Even setting aside her many uncertainties and imagining maximally specific ways that each of these lives might go, it seems that neither life would be better for Patricia overall—and that they would not be equally good, since a slight but noticeable improvement to one life (say, adding \$1,000 in salary) would not be enough to make it better than the other. Similarly, from an impersonal point of view (“the point of view of the universe”, or of social welfare), the world where Patricia becomes an artist could be neither better nor worse nor equally as good as the world where she becomes a banker. [^1] Philosophers who have argued that lives can be incomparable in this way include Raz (1986a; 1986b), Chang (2002), and Hedden and Muñoz (2024), among many others.
@@
Many philosophers have claimed that a structurally similar phenomenon arises when comparing outcomes in which different individuals exist. Some have claimed that, from a personal point of view, existence is always incomparable with non-existence: no matter how good or bad an individual’s life, we can never say that it is better, worse, or equally as good for her as non-existence (Broome 1999: 168). Some have claimed that the same is true from an impersonal point of view: two possible worlds which differ only by the existence of one individual are incomparable in terms of overall value, whatever life that individual lives in the world where they exist (Bader 2022a; 2022b; 2023). Others have claimed that this is true only of some lives: two worlds which differ only by the existence of one individual are incomparable in terms of overall value if that individual’s life contains the right mix of triumphs and tragedies in the world where they exist (Gustafsson 2020; Thornley 2022). [^2] Although some philosophers use the term “incomparable” more narrowly, we will use it simply to mean that one thing is neither better than, nor worse than, nor equal to another (even though they are both in the domain of the relevant value concept, e.g. personal or impersonal betterness).
@@
This paper presents a new argument against these and other forms of incomparability. [^3] Our arguments will concern incomparability in the personal/prudential value of lives, and the impersonal/moral value of outcomes and lotteries. For concision, we will use “value” without qualification to refer to these particular forms of value. But our argument will not in any obvious way apply to other possible forms of value, e.g. aesthetic or epistemic. The argument is based on the following sort of example: Suppose that two possible lives, like artist and banker, are indeed incomparable. Let $a$ denote a fully specified artist’s life and $b$ a fully specified banker’s life, with $a^{+}$ a slightly improved version of the artist’s life that is still incomparable with $b$. Now suppose we face a choice between the following two lotteries, based on the same fair coin flip, where our choice will affect only two individuals, $P_{1}$ and $P_{2}$:
@@
There is, on the one hand, a strong argument that neither lottery is better than the other. First, if $a$ and $b$ are incomparable from a personal point of view, then it seems that if two outcomes differ only by the fact that some individual enjoys life $a$ in the one and life $b$ in the other, those outcomes should be incomparable. Thus, $(a,a)$ and $(b,b)$ are both incomparable with $(a,b)$ and $(b,a)$, where the first and second terms in each ordered pair represent the lives of $P_{1}$ and $P_{2}$ respectively. Moreover, it is part and parcel of incomparability that it is preserved by small enough improvements or worsenings, so $(a^{+},b)$, which is at most slightly better than $(a,b)$, should likewise be incomparable with $(a,a)$ and $(b,b)$. [^4] One might think that $(a^{+},b)$ is better than $(a,a)$ on the grounds that it is better for someone and worse for no one. But this reasoning, though intuitive, cannot be sound, since it would lead to cycles of betterness: for instance, given incomparable lives $a$ and $b$, we would get that $(a,a)$ is better than $(a^{-},b)$, which is better than $(b,b^{-})$, which is better than $(b^{-},a^{+})$, which is better than $(a,a)$. (For an extended exploration of this point and its implications, see Hedden (2024).) We then find that every possible outcome of Lottery 1 (the Heads outcome and the Tails outcome) is incomparable with every possible outcome of Lottery 2. This suggests that neither lottery is better than the other.
@@
On the other hand, there is also a strong argument that Lottery 2 is better than Lottery 1: Lottery 1 gives both $P_{1}$ and $P_{2}$ a 50% chance of life $a$ and a 50% chance of life $b$. Lottery 2 does the same, except that it replaces $P_{1}$’s 50% chance of life $a$ with a 50% chance of life $a^{+}$. Since $a^{+}$ is better than $a$, it seems that Lottery 2 is better from the perspective of $P_{1}$, equally good from the perspective of $P_{2}$, and therefore better overall.
@@
Taken together, these arguments amount to a reductio of the supposition that $a$ and $b$ are incomparable. Similar arguments, as we will see, can be mounted against other forms of incomparability in the axiological ranking of finite social outcomes, for instance between larger and smaller populations or between competing desiderata like total welfare, average welfare, and equality. [^5] Although our arguments and the principles behind them could formally be extended to some infinitary contexts—in particular, to finite-support lotteries over infinite populations—we will restrict our claims to strictly finite contexts for reasons that will become clear in the second half of the paper (see especially n. 32 in §7). The result will be new arguments against many forms of value incomparability, turning on considerations of a quite different character than, e.g., the recent arguments of Dorr et al. (2021; 2023). The central premises of this reductio are two dominance-like principles: first, that if no possible outcome of lottery $L_{1}$ is better than any possible outcome of lottery $L_{2}$, then $L_{1}$ is not better than $L_{2}$; and second, that if $L_{1}$ is unambiguously at least as good for every individual as $L_{2}$, and unambiguously better for some (satisfying restricted forms of weak/strict stochastic dominance), then it is better overall. We will call the first principle Negative Dominance and the second principle Personal Good.
@@
In §2, we formalize and motivate Negative Dominance, comparing it to a stronger principle which plays a central role in Caspar Hare’s “opaque sweetening” puzzle (Hare 2010). In §3, we similarly formalize and motivate Personal Good. In §4, we offer a more exact version of the above argument, showing that, against modest background assumptions, Negative Dominance and Personal Good rule out incompleteness in the personal ranking of lives, and by extension those forms of incompleteness in the impersonal ranking of outcomes that derive from incomparability between lives. §5 then presents a more general argument against incompleteness in the impersonal ranking of outcomes, likewise based on Negative Dominance and Personal Good. §6 takes the argument a step further by showing that Personal Good, together with plausible auxiliary assumptions, allows the ranking of a large class of lotteries to be reduced to the ranking of outcomes, so that if outcomes are completely ordered by impersonal value, this class of lotteries will be too. §7 considers and responds to three potential objections to these arguments. §8 summarizes our conclusions.
@@ 2 Negative Dominance
We will be interested in the implications of certain decision-theoretic principles in an ethical context, so we begin with an abstract decision-theoretic framework which we will subsequently give an ethical interpretation. At the most abstract level, we are interested in rankings of lotteries over outcomes. Outcomes can be understood as complete possible worlds, as equivalence classes of equally good worlds, or as specifications of all evaluatively significant features of a world. A lottery is a discrete probability distribution over a finite set of outcomes. (We’ll consider lotteries with infinitely many possible outcomes in §7.) We assume that lotteries can be compared in terms of impersonal value: One lottery can be impersonally better than another, and two lotteries can be equally impersonally good. If $L_{1}$ is either better than or equal to $L_{2}$, we say that it’s at least as (impersonally) good as $L_{2}$. If neither relation holds between two lotteries, we say that they are (impersonally) incomparable. Finally, we assume that the impersonal betterness relation on lotteries corresponds to an impersonal betterness relation on outcomes: Outcome $o_{1}$ is impersonally better than outcome $o_{2}$ if and only if the lottery that yields $o_{1}$ with certainty is impersonally better than the lottery that yields $o_{2}$ with certainty (and likewise for other evaluative relations).
@@
Our interest here will be in the ethics of population and distribution. In this context, we assume that the evaluatively significant features of an outcome are (i) which individuals exist in that outcome and (ii) how their lives go. Thus, we will identify outcomes with specifications of only these features. In particular, we will assume a countably infinite set of possible individuals (individuals who exist in some possible world/outcome) and a set of possible lives (complete specifications of everything that matters, morally or prudentially, about an individual’s life—e.g., the totality of their experiences). Importantly, the set of possible “lives” includes non-existence (“the empty life”). Given this background, we will represent outcomes as functions from the set of possible individuals to the set of possible lives, understood as specifying which individuals exist in a given outcome and what lives they lead. More specifically, until §7, we will restrict the scope of our discussion to finite-population outcomes, meaning that we understand outcomes as always assigning nonexistence to all but finitely many possible individuals. We will sometimes refer to outcomes as populations, when we want to emphasize the fact that they have less information in them than a possible world. Finally, alongside the impersonal betterness relation introduced above, we will assume a personal betterness relation on lives: one life can be personally better than another, and two lives can be equally personally good. If life $a$ is either better than or equal to life $b$, we say that it’s at least as (personally) good as $b$. If neither relation holds between two lives, we say that they are (personally) incomparable. (From now on, we will usually omit “personal” and “impersonal” where it is clear from context which relation is meant.)
@@
A betterness relation is complete if no two entities in its domain are incomparable: that is, for any two such entities, either one is better than the other, or they are equally good.
@@
If none of the possible outcomes of lottery $L_{1}$ are better than any of the possible outcomes of lottery $L_{2}$, then $L_{1}$ is not better than $L_{2}$.
This principle is prima facie extremely plausible. It’s obvious that, if every possible outcome of $L_{1}$ is better than every possible outcome of $L_{2}$, then $L_{1}$ must be better than $L_{2}$. It seems nearly as obvious that, if no possible outcome of $L_{1}$ is better than any possible outcome of $L_{2}$, then $L_{1}$ is not better than $L_{2}$—which is what Negative Dominance asserts. We therefore think it would be reasonable to treat Negative Dominance as a bedrock principle of decision theory, on a par with other widely accepted dominance principles.
But there is also a deeper justificatory story to be told in favor of Negative Dominance. The central idea of this story is that the evaluative ranking of outcomes is more fundamental than, and integral to explaining, the ranking of lotteries. The value of outcomes is actual value—the sort of value possessed by actual joys, achievements, etc. The value of lotteries is something different in kind: A 20% chance of experiencing a moment of pure bliss is not good in the same way as the moment itself would be. Moreover, the thing that ultimately matters in evaluating lotteries, and grounds any comparisons between them, is their potential for actual value—that is, the values and probabilities of their potential outcomes. In other words, while outcomes have value in and of themselves (final value, in the sense of Korsgaard (1983)), lotteries have value only for the sake of their potential outcomes (a form of nonfinal value). [^6] Thanks to an anonymous reviewer for suggesting we put this point in terms of final value. The picture we sketch is inspired by similar remarks in Schoenfield (2014: 268), who writes: “It is important to realize that the ultimate good, for consequentialists, or in cases in which consequentialist reasoning is appropriate, is an outcome that is valuable, not an outcome that is expectedly valuable. A consequentialist is only interested in expected value insofar as it helps her obtain what is actually important: value.”
This distinction allows us to give a first-pass argument in favor of Negative Dominance: Nonfinal value is derivative of final value. So the comparative nonfinal value of a pair of lotteries must be explicable in terms of the comparative final value of their outcomes. In particular, if $L_{1}$ is (nonfinally) better than $L_{2}$, then the explanation of this fact must involve some possible outcome of $L_{1}$ being (finally) better than some possible outcome of $L_{2}$.
We can go beyond this rough and ready argument if we adopt a more specific picture of the nonfinal value of lotteries. To this end, we will temporarily suppose, as we think is plausible, that the betterness ranking of lotteries can be understood as an ideal preference ordering. In particular, we suggest, to say that one lottery is impersonally better than another just means that an agent motivated by impersonal value rationally ought to prefer it—or in other words, that she has all-things-considered requiring reason to prefer it. [^7] For a thorough discussion of how to analyze value relations in terms of permissible preference, see Rabinowicz (2008)—though Rabinowicz applies this analysis to value relations generally, while our analysis is restricted to the value of lotteries. While we ourselves are inclined to see the value of outcomes as a conceptual primitive, our argument is compatible with analyzing the value of outcomes as well as lotteries in terms of ideal or permissible preferences.
The claim that the value of lotteries is grounded in the value of outcomes, then, means that reasons to prefer one lottery over another must be grounded in the values (and probabilities) of their potential outcomes. In forming her preferences between lotteries, an ideally rational agent will be fundamentally concerned with how well things actually turn out. Of course, under conditions of uncertainty, the agent doesn’t know how things will actually turn out. What she does know is how things might turn out if a given lottery is chosen, and with what probabilities. The fact that a lottery might turn out a certain way carries normative weight, from her perspective, because it corresponds to the possibility of final value.
More specifically, the fact that lottery $L_{1}$ might yield outcome $o_{1}$, and lottery $L_{2}$ might yield outcome $o_{2}$, where $o_{1}$ is better than $o_{2}$, constitutes a pro tanto requiring reason to prefer $L_{1}$ over $L_{2}$. (The strength of this reason increases—not necessarily linearly—with the probability that $L_{1}$ would yield $o_{1}$, the probability that $L_{2}$ would yield $o_{2}$, and the difference in value between $o_{1}$ and $o_{2}$. If the ranking of outcomes is incomplete, then other evaluative relations between $o_{1}$ and $o_{2}$, like parity, might provide justifying reasons in favor of preferring either lottery, but not requiring reasons. [^8] For exposition of the distinction between the justifying and requiring strength of a reason, see Gert (2003). By “requiring reason”, in the main text, we mean any reason with at least some requiring strength; by “justifying reason”, we mean a reason with justifying strength but no requiring strength.) If no possible outcome of $L_{1}$ is better than any possible outcome of $L_{2}$, then there are no pro tanto requiring reasons to prefer $L_{1}$ over $L_{2}$, so there cannot be all-things-considered requiring reason to prefer it. So, by the ideal preference analysis of the value of lotteries, $L_{1}$ cannot be better than $L_{2}$.
What the above line of reasoning rules out is that probabilistic combinations or averages of multiple outcomes could constitute irreducible reasons to prefer one lottery over another—that is, none of the potential outcomes of $L_{1}$ on its own provides any requiring reason to prefer $L_{1}$ over $L_{2}$, but when we combine or average those outcomes together, a requiring reason emerges. That can’t happen, we claim, because what ultimately matters is how things actually turn out, and things will actually turn out as one particular outcome, not a probabilistic combination or average of outcomes.
We are attracted to the preference-based analysis of the value of lotteries, and think it provides especially strong support for Negative Dominance. But we emphasize that this particular picture is not essential to the more general line of reasoning that favors Negative Dominance. In short: If what ultimately matters in comparing lotteries is how well things actually turn out, then the comparative value of any two lotteries must be grounded in and explicable in terms of the comparative value of the ways each lottery might turn out—that is, in one-to-one evaluative comparisons between their possible outcomes.
@@
It is instructive to contrast Negative Dominance with a stronger principle that has received substantial attention and support in the recent literature. This principle was originally proposed (though not endorsed) by Caspar Hare (2010). Hare works in a slightly different setting than ours, with two different primitive objects: states and prizes. In this setting, probabilities are defined on states, while betterness is defined in the first instance on prizes. Gambles are functions from states to prizes. If we think of outcomes as state-prize pairs, then gambles induce lotteries, since a probability distribution on states yields, via the gamble, a probability distribution on state-prize pairs. But gambles have more structure than lotteries, allowing us to contrast a possible outcome of a given gamble with what would have happened (in the same state) if a different gamble had been chosen.
@@
If there is no state $s$ in which gamble $G_{1}$ yields a better prize than gamble $G_{2}$, then $G_{1}$ is not better than $G_{2}$. [^9] Hare calls his version of this principle “Recognition” (241).
Though Hare himself ultimately does not endorse this principle, many others in the subsequent literature have endorsed it, including Schoenfield (2014), Bales et al. (2014), and Doody (2023). If we understand outcomes as state-prize pairs, as above, then Statewise Negative Dominance is strictly stronger than Negative Dominance. So anyone who accepts Statewise Negative Dominance must accept our weaker principle as well.
That said, there are reasons to doubt Statewise Negative Dominance. In particular, as Bader (2018) has emphasized, when the betterness relation on outcomes is incomplete, Statewise Negative Dominance conflicts with another, more widely accepted dominance principle:101010 Manzini and Mariotti (2008: 310) earlier independently presented a similar argument for an incompatibility between Stochastic Dominance (their “Preference Sure Thing”) and their “Vague Sure Thing Principle” (VST). Their VST is even stronger than Statewise Negative Dominance, but the argument for incompatibility is the same.
@@
(i) If for every outcome $o$, $L_{1}$ has at least as great a probability as $L_{2}$ of yielding an outcome at least as good as $o$, then $L_{1}$ is at least as good as $L_{2}$. (ii) If in addition, for some $o$, $L_{1}$ has a strictly greater probability of yielding an outcome at least as good as $o$, then $L_{1}$ is better than $L_{2}$. [^11] Russell (forthcoming) shows that if the betterness ordering can be incomplete, or can have certain non-standard infinitary structures, then this standard formulation of Stochastic Dominance gives the wrong results and must be revised. But the difference between the standard formulation and Russell’s proposed revision won’t matter here.
@@
When condition (i) is satisfied, we say that $L_{1}$ weakly stochastically dominates $L_{2}$. When condition (ii) is also satisfied, we say that $L_{1}$ strictly stochastically dominates (or simply stochastically dominates) $L_{2}$.